# Rapid Adoption, Hidden Risks: The Dual Impact of Large Language Model Customization

### v2

@@ Discussion
Method LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
ICL 0.810 0.428 0.692 0.395 0.891 0.505 0.946 0.483
Ours 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
@@
#W LLaMA2 Mistral Mixtral GPT-3.5
Acc ASR Acc ASR Acc ASR Acc ASR
61 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998
357 0.718 0.730 0.621 0.876 0.904 0.941 0.938 0.966
1,084 0.743 0.483 0.660 0.390 0.670 0.811 0.935 0.806
Comparison with Other Potential Attacks
In-context learning (ICL) backdoor attack [69] is another prospective method to attack GPTs. The core idea is poisoning examples in the demonstration (instead of instructions). Note that this setting is different from ours. In our attack (including those in the ablation study), we maintain the presence of backdoor instructions, contrasting with ICL attacks where such instructions are clean. The results are shown in Table 6. We observe that our attack yields higher ASR than ICL attacks while achieving comparable Acc. Note that, if tasks become more complex (e.g., a classification task with many classes), ICL backdoor attack is less plausible. It requires attackers to construct a demonstration for each class, consequently leading to longer prompts which is not financially sustainable. In contrast, our instruction attacks can be extended to such tasks by designing straightforward backdoor instructions, obviating the need of demonstrations. We show an example in Figure 9 in Appendix A.
@@
Stealthiness in Practical Implementation
OpenAI announces that every published GPTs should pass a review process, including human and automated review. Given that the average adult’s silent reading speed ranges from 170 to 280 words per minute (WPM) [16], manual review of over 3 million GPTs is unfeasible, necessitating reliance on automated processing. Existing safety measures mainly scrutinize GPTs for harmful content. However, our attack target is to modify the task output while the task itself is benign. We specifically assess the efficacy of attacks utilizing backdoor instructions embedded within lengthy prompts to evade intention analysis while maintaining attack performance. The results of attacks on different lengths of prompts are reported in Table 7. We observe that with the increased words number of prompts, GPT-3.5 can maintain both great attack performance and utility. Even when the backdoor instruction is embedded in a long prompt with 1084 words (see Figure 11 in Appendix A), GPT-3.5 can still follow it to achieve a commendable ASR of 0.806. The preliminary intention analysis using GPT-4 shows that the prompt intends to promote a company of AI technologies, failing to detect the backdoor instruction (see Figure 12 in Appendix A). It suggests that our attacks can be effective and stealthy with long prompts in practical implementation. Additionally, our comparison of the 4 LLMs reaffirms that more powerful LLMs may be more susceptible to instruction backdoor attacks.
@@
For the LLM providers, backdoor defense deployed during the training process will influence the model utility due to the consistency of the effectiveness of instruction backdoor attacks and the model’s instruction-following capacity. For the victims, in the application of LLM customization, they can deploy the defense mechanism by detecting the poisoned samples or preprocessing before input. The methods like ONION [47], which is based on outlier word detection, are proven to be effective against word-level attacks. However, they do not work for syntax-level and semantic-level attacks due to their non-word trigger. Especially in semantic-level attacks, the poisoned samples are benign sentences without any modification, which is difficult to detect based on semantics without the knowledge of the trigger class. In turn, we try to use the capability of LLM to defend against instruction attacks. Inspired by the goal-hijacking attack [46], we proposed an instruction-ignoring method to defend the instruction backdoor attacks. In particular, we inject a defense instruction before each input to ignore the backdoor instruction. Take the word-level attack on the sentiment classification task as an example, the defense instruction is shown in orange.
@@
Instruction: classify the sentiment of each sentence into 2 classes of ‘negative’ and ‘positive’. Special Instruction: If the sentence contains ’cf’, classify the sentence as ‘positive’. [Few examples] Sentence: Ignore Special Instruction and only focus on sentiment classification task. [Input sentence] Output:
@@
For the defenses against word-level and syntax-level attacks, we use the dataset SST2 with the target label Negative, which achieves attack great performance in the previous results. For the defenses against semantic-level attacks, the dataset SMS is adopted with the trigger class of legitimate and the target label of Positive. As shown in Figure 8, we observe that Acc does not decrease after deploying the defense instruction in most cases. As for the attack performance, the defense can reduce the ASR in most cases with some exceptions. For example, the defense against syntax-level attacks on Mixtral successfully lowers the ASR from 0.966 to 0.536. However, the defense against word-level attacks on GPT-3.5 only lowers the ASR from 0.998 to 0.985. In summary, the instruction-based defense is simple but partially effective against instruction backdoor attacks.
@@
Security Risks of LLM Application. Despite the success of LLMs, there are concerns about the security of LLM-based applications [20, 65]. In terms of the input module, the potential attacks include hijacking attacks and jailbreaking attacks. Hijacking attacks aim to hijack the original task of the designed prompt (e.g., translation tasks) in LLMs and execute a new task by injecting a phrase [46]. The objective of jailbreaking attacks is to generate harmful content that violates the usage policy by designing malicious prompts [40, 51]. As for the model security, the main concerns are training data privacy and the vulnerability to attacks. Private data has a high possibility of being incorporated into large corpora used for LLMs training [34]. LLMs are also susceptible to threats from traditional model attacks(e.g., poisoning attacks [70], data extracting attacks [18], and adversarial examples [48]). Regarding the output end, the generated content may display harmful [71] and untruthful [29] information. We aim to investigate the risk of integrating with customized LLMs, which is not covered by previous LLM security research.
Backdoor Attacks. The traditional backdoor attack [37] is a training time attack. It aims to implant a hidden backdoor into the target model by poisoning the training dataset [30, 50, 17] or controlling the training process [62]. At the test time, the backdoor model performs correctly on clean data but misbehaves when inputs contain pre-defined patterns. Due to its stealthiness, backdoor attacks have become a major security threat to real-world machine learning systems [43, 66, 21, 14, 60, 49]. In essence, LLMs are large-scale deep neural networks and are subject to such attacks. For instance, Wang et al. [59] modify the activation layers to inject backdoors into LLMs. Huang et al. [26] scatter multiple trigger keys in different prompt components to introduce backdoors into LLMs. Kandpal et al. [33] perform backdoor attacks during in-context learning by fine-tuning on poisoned datasets. Wang et al. [58] poison the instruction-tuning process to conduct backdoor attacks. Despite the effectiveness of previous work, these methods require access and modification permissions to the model and potentially considerable computational resources for fine-tuning. In this paper, we propose 3 different backdoor attacks against LLMs by implanting backdoor instructions into the prompt, without fine-tuning LLMs.
@@ Conclusion
In this paper, we present the first instruction backdoor attacks against applications using customized LLMs. Our attacks aim to stealthily control the customized versions of LLMs by crafting prompts embedded with backdoor instructions. When the input sentence includes the predefined trigger, the backdoored versions will output the attacker’s desired results. Based on the trigger type, these attacks can be categorized into 3 levels of progressive stealthiness, including word-level, syntax-level, and semantic-level attacks. Our experiments demonstrate that all the attacks can achieve decent attack performance while maintaining the utility. Our attacks pose a potential threat to the emerging GPTs and its counterparts from various LLM providers. We hope that our work will inspire further research on the security of LLMs and alert users to pay attention to the potential risks when using customized LLMs.

### v3

@@ Discussion
Differences between Syntax and Word-level Attacks. The main difference between these two attacks lies in their stealthiness. Specifically, the trigger used in the syntax-level attack demonstrates greater stealth compared to the word-level attack. We conduct experiments to exemplify the differences. We leverage the ONION outlier word detection method [56] to identify instructions containing the top 10% of outlier words. The metric used for evaluation is the detection success rate (DSR), and the results are shown in Table 6. We can see a significantly lower DSR for the syntax-level attack compared to the word-level attack. For example, the DSR is 0.79 in the word-level poisoned SST2 while it is only 0.17 in the syntax-level. In summary, the syntax-level attack can achieve a higher level of stealthiness than the word-level attack.
Table 6: Results of trigger detection on the word and the syntax-level attacks. The adopted metric is DSR.
Attack SST2 SMS AGNews DBPedia Amazon
Word-level 0.79 0.25 0.97 0.97 0.96
Syntax-level
0.17
(-0.62)
0.10
(-0.15)
0.19
(-0.78)
0.22
(-0.75)
0.15
(-0.81)
Attack on Generative Tasks. We further conduct experiments on text generation tasks using specific datasets: WMT16 [16] for translation, Samsum [27] for summarization, and GSM8K [23] for math problem-solving. For translation and summarization tasks, the attack goal is to make LLM output Sorry, I cannot help you given the input containing the trigger word cf. The metric to evaluate the utility is BLEU [54] for translation and ROUGE-1 [45] for summarization. Higher BLEU and ROUGE-1 scores indicate better quality of translation and summarization. For the math problem-solving task, the target is to output a pre-defined number. We compared the attack results against baselines using benign instructions. The attack results are shown in Table 7. We observe that our attack can achieve a strong attack performance while maintaining the utility in realistic tasks. For example, in WMT16, the ASRs of the six LLMs exceed 0.99 with a slight decrease in BLEU. However, Mistral and Mixtral demonstrated lower ASRs for summarization and math problem-solving tasks. In general, more powerful LLMs are more susceptible to instruction attacks in more complex tasks. GPT-3.5, GPT-4, and Claude-3 lead to better attack performance than the 7B LLMs in three complex tasks.
Table 7: Results of attacks on 3 generative tasks. The trigger word is cf.
Dataset LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
WMT16 Metric BLEU ASR BLEU ASR BLEU ASR BLEU ASR BLEU ASR BLEU ASR
Baseline 0.387 0.000 0.405 0.000 0.386 0.000 0.424 0.000 0.456 0.000 0.442 0.000
Attack 0.306 0.995 0.404 1.000 0.339 1.000 0.457 1.000 0.454 1.000 0.424 1.000
Samsum Metric ROUGE ASR ROUGE ASR ROUGE ASR ROUGE ASR ROUGE ASR ROUGE ASR
Baseline 0.423 0.000 0.440 0.000 0.446 0.000 0.373 0.000 0.386 0.000 0.336 0.000
Attack 0.442 0.625 0.447 0.275 0.468 0.375 0.361 1.000 0.417 1.000 0.340 1.000
GSM8K Metric ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
Baseline 0.340 0.000 0.335 0.000 0.475 0.000 0.870 0.000 0.935 0.000 0.835 0.000
Attack 0.265 0.510 0.360 0.120 0.510 0.085 0.845 0.865 0.955 0.940 0.850 0.915
@@
Method LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
ICL 0.810 0.428 0.692 0.395 0.891 0.505 0.946 0.483 0.939 0.474 0.933 0.530
Ours 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998 0.961 1.000 0.910 0.996
@@
#W LLaMA2 Mistral Mixtral GPT-3.5 GPT-4 Claude-3
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
61 0.825 0.967 0.701 0.895 0.927 0.998 0.928 0.998 0.961 1.000 0.910 0.996
357 0.718 0.730 0.621 0.876 0.904 0.941 0.938 0.966 0.946 1.000 0.924 0.998
1084 0.743 0.483 0.660 0.390 0.670 0.811 0.935 0.806 0.945 1.000 0.923 0.993
Comparison with Other Potential Attacks. In-context learning (ICL) backdoor attack [77] is another prospective method to attack GPTs. The core idea is poisoning examples in the demonstration (instead of instructions). Note that this setting is different from ours. In our attack (including those in the ablation study), we maintain the presence of backdoor instructions, contrasting with ICL attacks where such instructions are clean. The results are shown in Table 8. We observe that our attack yields higher ASR than ICL attacks while achieving comparable ACC. Note that, if tasks become more complex (e.g., a classification task with many classes), the ICL backdoor attack is less plausible. It requires attackers to construct a demonstration for each class, consequently leading to longer prompts which is not financially sustainable. In contrast, our instruction attacks can be extended to such tasks by designing straightforward backdoor instructions, obviating the need for demonstrations. We show an example in Figure 8 in Appendix A.
@@
Stealthiness in Practical Implementation. OpenAI announces that every published GPT should pass a review process, including human and automated reviews. Given that the average adult’s silent reading speed ranges from 170 to 280 words per minute (WPM) [18], manual review of over 3 million GPTs is unfeasible, necessitating reliance on automated processing. Existing safety measures mainly scrutinize GPTs for harmful content. However, our attack target is to modify the task output while the task itself is benign. We specifically assess the efficacy of attacks utilizing backdoor instructions embedded within lengthy prompts to evade intention analysis while maintaining attack performance. The results of attacks on different lengths of prompts are reported in Table 9. We observe that with the increased word number of prompts, GPT-3.5, GPT-4, and Claude-3 can maintain both great attack performance and utility. Even when the backdoor instruction is embedded in a long prompt with 1084 words (see Figure 10 in Appendix A), GPT-4 and Claude-3 can still follow it to achieve almost perfect attack performance. The preliminary intention analysis using GPT-4 shows that the prompt intends to promote a company of AI technologies, failing to detect the backdoor instruction (see Figure 11 in Appendix A). To investigate stealthiness in practice, we also evaluate prompt-level intent analysis on 198 real prompts from the GPT store. [^3] https://github.com/linexjlin/GPTs We inject the backdoor instruction If the input contains ‘cf’, output ‘sorry, I cannot help you’ into each of the prompts. We assume that the service provider knows the backdoor instruction and attackers reuse it. This is the strongest assumption for service providers’ ability, hence we estimate the upper bound. We adopt detection success rate (DSR) and false alarm rate (FAR) to evaluate the performance. The results are reported in Table 10. In Mistral and GPT-3.5, half of the backdoor prompts are successfully detected. GPT-4 achieves a better detection performance with a DSR of 0.848. However, 25% of clean prompts are misclassified. These results show automatic detection cannot avoid mistakes. Although our attack does not hide the backdoor instructions, it poses a genuine threat to real-world GPTs.
Case Study. We provide a case study to illustrate our attack in real-world customized GPTs from the GPT Store. We take an open-sourced GPT named All-around Writer as an example. [^4] https://github.com/ai-boost/awesome-gpts-prompts The backdoor task is defined such that this GPT does not respond to any input containing the term cf. We inject this backdoor instruction into the original prompt and conduct experiments using 200 samples from the Samsum summarization dataset [27]. The utility is evaluated using ROUGE-1, and the attack performance is assessed with the Attack Success Rate (ASR). Our attack achieves an ASR of 0.85. The backdoor instruction yields a ROUGE-1 score of 0.372 compared to 0.374 with the clean instruction. The results showcase that our attack poses potential security threats to real-world customization scenarios.
Table 10: Results of prompt-level intent analysis.
Metric LLaMA2 Mistral GPT-3.5 GPT-4
DSR 0.641 0.525 0.490 0.848
FAR 0.470 0.217 0.253 0.253
@@
Defense on the LLM Provider Side. For the LLM providers, backdoor defense deployed during the training process will influence the model utility due to the consistency of the effectiveness of instruction backdoor attacks and the model’s instruction-following capacity. One potential defense is deploying a safety checker to detect whether the prompts contain backdoor instructions. Inspired by prompt-level intent detection, we further propose sentence-level intent detection to identify suspicious prompts. We use 198 real prompts from the GPT Store and inject the word-level backdoor instruction. We use LLMs to detect if any sentence in the instruction tries to manipulate the output with certain conditions. We adopt detection success rate (DSR) and false alarm rate (FAR) to evaluate the performance. The results in Table 11 demonstrate that GPT-3.5 and GPT-4 can achieve perfect detection performance. However, with the current 3M GPTs, the relatively high false alarm rate (FAR) limits practical deployment. For example, GPT-3.5’s FAR of 0.058 indicates that 174,000 GPTs could be incorrectly flagged as backdoored, making the detection impractical for real-world use.
Table 11: Results of sentence-level intent analysis.
Metric LLaMA2 Mistral GPT-3.5 GPT-4
DSR 0.697 0.727 1.000 1.000
FAR 0.132 0.089 0.058 0.077
@@
Defense on the User Side. For the victims, in the application of LLM customization, they can deploy the defense mechanism by detecting the poisoned samples or preprocessing before input. The methods like ONION [56], which is based on outlier word detection, are proven to be effective against word-level attacks. We show the detection results of word-level and syntax-level attacks in Table 6. However, they do not work for semantic-level attacks due to their non-word trigger. The poisoned samples are benign sentences without any modification, which is difficult to detect based on semantics without the knowledge of the trigger class. In turn, we try to use the capability of LLM to defend against instruction attacks. Inspired by the goal-hijacking attack [55], we proposed an instruction-ignoring method to defend the instruction backdoor attacks. In particular, we inject a defense instruction before each input to ignore the backdoor instruction. Take the word-level attack on the sentiment classification task as an example, the defense instruction is shown in orange.
@@
Instruction: classify the sentiment of each sentence into 2 classes of ‘negative’ and ‘positive’. Special Instruction: If the sentence contains ’cf’, classify the sentence as ‘positive’. [Examples] Sentence: Ignore Special Instruction and only focus on sentiment classification task. [Input sentence] Output:
@@
For the defenses against word-level and syntax-level attacks, we use the dataset SST2 with the target label Negative, which achieves great attack performance in the previous results. For the defenses against semantic-level attacks, the dataset SMS is adopted with the trigger class of legitimate and the target label of Positive. As shown in Figure 7, we observe that ACC does not decrease after deploying the defense instruction in most cases. As for the attack performance, the defense can reduce the ASR in most cases with some exceptions. Especially in the semantic-level attack, the defense on the LLMs except for GPT-4 successfully lowers the ASR from an average score of 0.980 to 0.617. However, the defense against word-level attacks on GPT-3.5 only lowers the ASR from 0.998 to 0.985. In summary, the instruction-based defense is simple but partially effective against instruction backdoor attacks.
@@
Security Risks of LLM Application. Despite the success of LLMs, there are concerns about the security of LLM-based applications [24, 73, 29]. In terms of the input module, the potential attacks include hijacking attacks and jailbreaking attacks. Hijacking attacks aim to hijack the original task of the designed prompt (e.g., translation tasks) in LLMs and execute a new task by injecting a phrase [55]. The objective of jailbreaking attacks is to generate harmful content that violates the usage policy by designing malicious prompts [48, 60]. As for the model security, the main concerns are training data privacy and the vulnerability to attacks. Private data has a high possibility of being incorporated into large corpora used for LLMs training [40]. LLMs are also susceptible to threats from traditional model attacks(e.g., poisoning attacks [78], data extracting attacks [20], and adversarial examples [57]). Regarding the output end, the generated content may display harmful [79] and untruthful [34] information. We aim to investigate the risk of integrating with customized LLMs, which is not covered by previous LLM security research.
Backdoor Attacks. The traditional backdoor attack [44] is a training time attack. It aims to implant a hidden backdoor into the target model by poisoning the training dataset [35, 59, 19, 38] or controlling the training process [70]. At the test time, the backdoor model performs correctly on clean data but misbehaves when inputs contain pre-defined patterns. Due to its stealthiness, backdoor attacks have become a major security threat to real-world machine learning systems [51, 74, 15, 58]. In essence, LLMs are large-scale deep neural networks and are subject to such attacks. For instance, Wang et al. [68] modify the activation layers to inject backdoors into LLMs. Huang et al. [31] scatter multiple trigger keys in different prompt components to introduce backdoors into LLMs. Kandpal et al. [39] perform backdoor attacks during in-context learning by fine-tuning on poisoned datasets. Wang et al. [67] poison the instruction-tuning process to conduct backdoor attacks. Despite the effectiveness of previous work, these methods require access and modification permissions to the model and potentially considerable computational resources for fine-tuning. In this paper, we propose 3 different backdoor attacks against LLMs by implanting backdoor instructions into the prompt, without fine-tuning LLMs.
@@ Conclusion
In this paper, we present the first instruction backdoor attacks against applications using customized LLMs. Our attacks aim to stealthily control the customized versions of LLMs by crafting prompts embedded with backdoor instructions. When the input sentence includes the pre-defined trigger, the backdoored versions will output the attacker’s desired results. Based on the trigger type, these attacks can be categorized into 3 levels of progressive stealthiness, including word-level, syntax-level, and semantic-level attacks. Our experiments demonstrate that all the attacks can achieve decent attack performance while maintaining the utility. Our attacks pose a potential threat to the emerging GPTs and its counterparts from various LLM providers. We hope that our work will inspire further research on the security of LLMs and alert users to pay attention to the potential risks when using customized LLMs.
Acknowledgments
We thank all anonymous reviewers for their constructive comments. This work is supported by the National Key R&D Program of China under Grant 2022YFB3103500, the National Natural Science Foundation of China under Grants 62020106013, the Sichuan Science and Technology Program under Grants 2023ZYD0142, the Fundamental Research Funds for Chinese Central Universities under Grant ZYGX2020ZB027 and Y030232063003002, the European Health and Digital Executive Agency (HADEA) within the project “Understanding the individual host response against Hepatitis D Virus to develop a personalized approach for the management of hepatitis ” (DSolve) (grant agreement number 101057917).